the average of three consecutive numbers is 26. What are the three numbers?
Found 2 solutions by Cromlix, Alan3354:
Answer by Cromlix(4381)

Three consecutive numbers:
n, n + 1, n + 2
Average of these three numbers = 26
(n + n + 1 + n + 2)/3 = 26
Cross multiply:
(n + n + 1 + n + 2)= 78
Collect like terms:
3n + 3 = 78
3n = 75
n = 25
Three consecutive numbers:
25, 26, 27.
Hope this helps.
;-)

Answer by Alan3354(69443)

the average of three consecutive numbers is 26. What are the three numbers?
------------
The average is the middle number --> 25, 26, 27
-------------
Or,
(n-1) + n + (n+1) = 26*3
3n = 78
n = 26